# Integrated Communication, Navigation, and Remote Sensing in LEO Networks with Vehicular Applications

### v1

@@ IV A Fundamental Case Study
This section evaluates performance gain achieved by the lowest two levels of integration. For the traditional situation without integration, denote the numbers of the CNR satellites to be $M_{c}$, $M_{n}$, and $M_{r}$, respectively, and the available spectra for the CNR satellites to be $B_{c}$, $B_{n}$, and $B_{r}$, respectively. Each type of the satellites is uniformly distributed on the sphere with an altitude of 500 km. In the case of function-level integration, there are equivalent $M_{c}+M_{n}+M_{r}$ satellites uniformly distributed on the same sphere, each of which has the CNR functionalities simultaneously. One step further, we assume ideal waveform design such that all the CNR functionalities can share the $B_{c}+B_{n}+B_{r}$ bandwidth without performance tradeoffs in the signal-level integration. There are 1,400 vehicular users distributed on circular cap area of $7.36\times 10^{5}\ {\rm km}^{2}$ near the equator. Each user has a minimum required data rate for communication and a data rate lower than this threshold causes service outage. Meanwhile, every user needs positioning, timing, and remote sensing information downloading, where the LEO enhanced positioning is performed on the basis of the GPS. In the integration cases, we consider each swath is illuminated by two satellites simultaneously to get observation resolution improvement, say, by a half, in the range direction by data fusion. The AoI of the remote sensing data received by the ground users is derived by adding satellite revisit time and the data downloading time. Some key simulation parameters are summarized in Table I.
@@
The overall performance of the CNR is presented by a spider web shown in Fig. 4, where the communication is evaluated by capacity outage probability and mean ergodic capacity marked in red, the navigation is characterized by positioning accuracy and timing error marked in blue, and the remote sensing is captured by resolution in range direction and AoI measured at the ground receivers marked in purple. It can be observed that both function-level and signal-level integration can comprehensively improve the CNR performance. The function-level integration introduces more available satellites for the three functionalities, which shortens the communications distance, provides more anchor location references for navigation, and decreases the revisit time for remote sensing. The gain of the signal-level integration mainly comes from expansion of the available spectrum for each functionality, which results in less improvement in communication regime as the increased amount of spectrum is smaller than other two functionalities. More importantly, the almost centimeter-level positioning achieved by the signal-level integration is quite fascinating for autonomous vehicles to perform lane navigation. Meanwhile, the decimeter-level earth observation accuracy is helpful to distinguish different vehicles on the ground, which can be leveraged to conduct sophisticated road traffic analysis and transportation system optimization. Such quantitative improvement of the navigation and remote sensing performance by even the lowest two levels of integration, e.g., decreasing the positioning error and view resolution by one order of magnitude, shines light on the future ICNR industry while communication performance is also expected to be enhanced significantly by deeply exploiting the navigation and remote sensing data in higher levels of integration. Finally, it is safe to admit the promising potentials of the ICNR in comprehensively enhancing the CNR performance as the integration level goes up, especially when the total number of the LEO satellites grows further to tens of thousands, which is promising to offer distinct transportation observation, centimeter-level positioning, and timely downloading of huge remote sensing data for future vehicular applications.
V Key Tradeoffs of ICNR Management
Optimization of the ICNR system has to jointly take the three functionalities into account. However, due to distinct working mechanisms of the three functionalities from one another, many performance tradeoffs need to be addressed, especially for vehicular applications.
@@ V-A Constellation Design
As the constellation size or orbit altitude increases, the communication coverage probability usually increases first due to reducing coverage holes and then decreases due to rising interference if no advanced interference coordination technique is implemented. On the other hand, the GDOP in the navigation regime monotonically decreases with the constellation size and the orbit altitude, and around 600 satellites deployed in the orbits with an altitude of 500 km can achieve the same performance with the GPS system [11]. However, only linear decrease of the GDOP may be achieved by exponential increase of the number of satellites greater than about 600. Furthermore, in addition to fast varying satellite distribution, the mobility of vehicles makes terminals experience more complicated positioning performance influenced by the constellation. For the remote sensing functionality, the constellation size has a direct influence on the temporal resolution in terms of the revisit time but not on the spatial resolution, where the revisit time in strip-map mode is generally discretely decreased with the number of satellites. It is worth noting that on-demand spotlight mode may be more suitable for transportation remote sensing for vehicular applications, where urban area has a higher requirement than rural area. Since a mega-constellation of LEO satellites must be launched batch by batch lasting for several years or more, it is thus critical to optimize the evolution route of the error-growing constellation, so as to march in lockstep with the increasing demand on the CNR in autonomous driving.
@@ V-B Satellite Association
Associating a terminal to the nearest access point has long been the fundamental principle in wireless communication field to experience the lowest channel fading, implying that the nadir of a satellite on the ground gains the best channel quality. However, the thing becomes totally different for the navigation and remote sensing. According to (4), directly looking below with a view angle of $\varphi=0$ fails to distinguish any two objects with any distance apart. Therefore, a SAR satellite needs to take a side look, and a typical view angle of $\varphi=30^{\circ}$ at an orbit altitude of $800$ km could lead to a distance of around 373 km between the swath center and the nadir [11]. Although the ground range resolution gets better as the view angle becomes larger, a moderate view angle is usually adopted to avoid severe signal attenuation. As for the navigation, the GDOP becomes better if the satellites are more scattered throughout the sky, indicating that the navigation functionality prefers to use signals emitted from far satellites dispersed in all directions. In other words, the navigation terminals may need to discard some signals from nearby satellites although many satellites can be seen in the era of mega-constellations of LEO satellites. The distinct association preference among different functionalities, as shown in Fig. 5, causes great difficulty in multi-purpose signal design and beam control toward moving vehicles. Additionally, on the terminal side, vehicles need to dynamically conduct satellite selection for each functionally to adapt to satellite distribution and own location variations.
@@ V-C Resource Management
On-demand service providing could probably be one of the main features of the LEO mega-constellation based ICNR systems. Particularly, as a small number of satellites are deployed in traditional remote sensing, the strip-map mode is commonly implemented to obtain earth observation images with global coverage, usually causing long revisit time greater than a couple of days or more. By contrast, since every ground point is able to see many satellites in the LEO mega-constellations, the spotlight mode is more appropriate to be performed whenever there is a particular demand on real-time observation towards a region while the strip-map mode is still conducted by a number of satellites to offer regular earth observation. Therefore, how many satellites work in the strip-map mode and how many are reserved for sporadic services should be carefully designed. In addition, as the CNR functionalities share the limited resources, such as power, computing, cache, and spectrum, smart allocation of these heterogeneous resources plays a prominent role in enhancing overall performance, including how to balance communication and computing payload by appropriately processing the remote sensing data in orbit [15]. Since a growing number of intelligent vehicles persistently impose requirements on high-quality earth observation, navigation, and remote sensing data downloading, sequential resource management decisions have to be optimized in response to network environment variations, which is pretty challenging in the highly dynamic ICNR networks with fast moving satellites and vehicles.
@@ VI Conclusion
In this article, we have proposed the ICNR framework in LEO satellite networks with vehicular applications. After presenting the most fundamental principles of the CNR functionalities, we have classified the proposed ICNR networks in different integration levels. The function-level and signal-level integration have then been instantiated as a numerical case study to show the integration gain. Finally, several key tradeoffs in managing the ICNR networks have been emphasized from the perspective of vehicular applications.

### v2

@@ IV A Fundamental Case Study
This section evaluates performance gain achieved by the lowest two levels of integration for autonomous driving in rural area without terrestrial network coverage. Each type of satellites is uniformly distributed on the sphere with an altitude of 500 km. In the case of function-level integration, the number of the ICNR satellites is set as the sum of the numbers of the traditional CNR satellites. One step further, we assume ideal waveform design such that all the CNR functionalities can share the overall bandwidth of the traditional CNR satellites without performance tradeoffs in the signal-level integration. There are 1,400 vehicles sparsely distributed on circular cap area of $7.36\times 10^{5}\ {\rm km}^{2}$, moving with a speed of less than 108 km/h. Each vehicle reports its coarse sensing information and driving maneuver intention to LEO satellites once every 3 meters, such that the ITS center can monitor the driving performance and also provide cooperative perception service to others. With the perception parameters adopted in [15], such periodical message delivery roughly requires a safeguarded data rate threshold 0.7 Mbps, and communication outage occurs if achievable data rate is lower than the threshold. Besides such safety-related message transfer, vehicles also requires best-effort downloading service to provide passengers with online infotainment, which cares more about the metric of ergodic capacity. Meanwhile, every user needs positioning, timing, and remote sensing information, where the LEO enhanced positioning is performed on the basis of the GPS. In the integration cases, we consider each swath is illuminated by two satellites simultaneously to get observation resolution improvement, say, by a half, in the range direction by data fusion. The AoI of the remote sensing data received by the ground users is derived by adding satellite revisit time and the data downloading time. Some key simulation parameters are summarized in Table I.
@@
The overall performance of the CNR is presented by a spider web shown in Fig. 4, where the communication is evaluated by capacity outage probability and mean ergodic capacity marked in red, the navigation is characterized by positioning accuracy and timing error marked in blue, and the remote sensing is captured by resolution in range direction and AoI measured at the ground receivers marked in purple. It can be observed that both function-level and signal-level integration can comprehensively improve the CNR performance. The function-level integration introduces more available satellites for the three functionalities, which shortens the communication distance, provides more anchor location references for navigation, and decreases the revisit time for remote sensing. The gain of the signal-level integration mainly comes from expansion of the available spectrum for each functionality, which results in less improvement in communication regime as the increased amount of spectrum is smaller than the other two functionalities. More importantly, near centimeter-level positioning achieved by the signal-level integration is quite fascinating for autonomous vehicles to perform lane navigation. Meanwhile, the decimeter-level earth observation accuracy is helpful to distinguish different vehicles on the ground, which can be leveraged to conduct sophisticated road traffic analysis and transportation system optimization. Such quantitative improvement of the navigation and remote sensing performance by even the lowest two levels of integration, e.g., decreasing the positioning error and discernible distance in view by one order of magnitude, shines light on the future ICNR industry, while communication performance is also expected to be enhanced significantly by deeply exploiting the navigation and remote sensing data in higher levels of integration. Finally, it is safe to admit the promising potentials of the ICNR in comprehensively enhancing the CNR performance as the integration degree goes up to data-level and application-level, by deep inter-functionality knowledge exploitation and application’s QoS requirement domain exploration.
V Future Research Directions
Optimization of the ICNR system has to jointly take the three functionalities into account. However, due to distinct working mechanisms of the three functionalities from one another, there are some trending challenges and open research directions in the ICNR framework, especially for vehicular applications.
@@ V-A Constellation Design
As the constellation size increases, the quality of communication may increase first due to shrinking coverage holes, and then decrease because of rising inter-satellite interference if no advanced interference coordination technique is implemented. On the other hand, with the LEO constellation getting dense, the GDOP in the navigation regime greatly decreases to around 2 in the early stage, while less benefit is anticipated by deploying more satellites thereafter [11]. Furthermore, in addition to fast varying satellite distribution, the mobility of vehicles makes terminals experience more complicated positioning performance influenced by the constellation. For the remote sensing functionality, the constellation size has a direct influence on the temporal resolution in terms of revisit time, where the revisit time in stripmap mode is generally decreasing with the number of satellites. It is worth noting that on-demand spotlight mode may be more suitable for transportation remote sensing for vehicular applications, where urban area has a higher requirement than rural area. Since a mega-constellation of LEO satellites must be launched batch by batch lasting for several years or more, it is thus critical to optimize the evolution route of the ever-growing constellation, so as to march in lockstep with the increasing CNR demand of autonomous driving.
@@ V-B Satellite Association
Associating a terminal to the nearest access point has long been the fundamental principle in wireless communication field to experience the lowest channel fading, implying that the nadir of a satellite on the ground gains the best channel quality. However, the thing becomes totally different for the navigation and remote sensing. In particular, directly looking below with a view angle of $0^{\circ}$ fails to distinguish any two objects with any distance apart in SAR based remote sensing. Therefore, a SAR satellite needs to take a side look, and a view angle of $25^{\circ}$ at an orbit altitude of $800$ km could lead to a distance of around 373 km between the swath center and the nadir. Although the ground range resolution gets better as the view angle becomes larger, a moderate view angle is usually adopted to avoid severe signal attenuation. As for the navigation, the GDOP becomes better if the satellites are more scattered throughout the sky, indicating that the navigation functionality prefers to use signals emitted from far satellites dispersed in all directions. In other words, the navigation terminals may need to discard some signals from nearby satellites although many satellites can be seen in the era of mega-constellations of LEO satellites. The distinct association preference among different functionalities, as shown in Fig. 5, causes great difficulty in multi-purpose waveform design and beam control towards moving vehicles in the proposed ICNR networks. Additionally, on the terminal side, vehicles need to dynamically conduct satellite selection for each functionality to adapt to satellite distribution and own location variations.
@@ V-C Resource Management
On-demand service providing could probably be one of the main features of the LEO mega-constellation based ICNR systems. Particularly, as a small number of satellites are deployed in traditional remote sensing, the stripmap mode is commonly implemented to obtain earth observation images with global coverage, usually causing long revisit time greater than a couple of days or more. By contrast, since every ground point is able to see many satellites in the LEO mega-constellations, the spotlight mode is more appropriate to be performed whenever there is a particular demand on real-time observation towards a region, while the stripmap mode can be still conducted by a number of satellites to offer regular earth observation. Therefore, how many satellites that work in the stripmap mode and how many that are reserved for sporadic services should be carefully designed. In addition, as the CNR functionalities share the limited resources, such as power, computing, cache, and spectrum, smart allocation of these heterogeneous resources plays a prominent role in enhancing overall performance, including how to balance communication and computing payload by appropriately processing the remote sensing data in orbit. Since a growing number of fast moving vehicles persistently impose comprehensive CNR requirements, sequential resource management decisions have to be optimized in response to user mobility and satellite topology variation, which is pretty challenging in the highly dynamic ICNR networks.
@@ VI Conclusion
In this article, we have proposed the ICNR framework in LEO satellite networks with vehicular applications. After presenting the most fundamental principles of the CNR functionalities, we have classified the proposed ICNR networks in different integration levels. The function-level and signal-level integration have then been instantiated as a numerical case study to show the integration gain. Finally, possible future research directions for more efficient exploitation of the ICNR framework have been highlighted.